# Sample Complexity of Distributionally Robust Average-Reward Reinforcement Learning

### v1

@@ 2 Literature Review
We provide a summary of state-of-the-art sample complexity results in the literature in Table 1.
Table 1: Summary of S.O.T.A. sample complexity results in the literature.
@@
Standard Discounted $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-3}\epsilon^{-2}})$ Azar et al., (2013); Li et al., (2020)
@@
DR-RL Discounted $\widetilde{O}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-4}\epsilon^{-2}})$ Shi and Chi, (2024); Wang et al., 2024c
@@ 3.2 Uniform Ergodicity
An MDP (or its controlled transition kernel $P$) is said to be uniformly ergodic if for all policies $\pi\in\Pi$, $t_{\mathrm{minorize}}:=\max_{\pi\in\Pi}t_{\mathrm{minorize}}(P_{\pi})<\infty$.
@@ 3.3 Distributionally Robust Discounted-Reward and Average-Reward MDPs
$\mathcal{T}_{\gamma}^{*}(V)(s):=$ $\max_{a\in\mathbf{A}}\left\{{r(s,a)+\gamma\Gamma_{\mathcal{P}_{s,a}}(V)}\right\}$ (3.3)
$\mathcal{T}^{*}(v)(s):=$ $\max_{a\in\mathbf{A}}\left\{{r(s,a)+\Gamma_{\mathcal{P}_{s,a}}(v)}\right\}$
@@
A DR-AMDP (or $\mathcal{P}$) is said to be uniformly ergodic if for all controlled kernels $Q\in\mathcal{P}$, $Q$ is uniformly ergodic as in Definition 3.3.
@@ 4.2 Anchored DR-AMDP
for a solution pair $(g_{\underline{\widehat{\mathcal{P}}}}^{*},v_{\underline{\widehat{\mathcal{P}}}}^{*})$
@@ 5 Numerical Experiments
This example clarifies our use of $m_{\vee}$ in Definition 3.3: while $m_{\pi}\equiv 1$ for all $\mathfrak{p}\in(0,1/2]$, the minorization time $t_{\mathrm{minorize}}$ is unbounded, approaching infinity as $\mathfrak{p}$ goes to 00.
@@
We then perform regression on data points on each MDP instance with the same parameter $\mathfrak{p}$. The plots demonstrate the error converging with rate $n^{-1/2}$, evidenced by the slope of $-1/2$ in Figure 2 on a log-log scale. We observe a remarkably low variance around the regression line of both algorithms, given that each data point is a single independent run of the corresponding algorithm.
@@ Appendix A Notations and Basic Properties
$\mathrm{Span}\left(u\right):=$ $\max_{1\leq i\leq d}(u_{i})-\min_{1\leq i\leq d}(u_{i})$ (A.1)
@@ Appendix B Uniform Ergodicity of the KL Uncertainty Set
>$\stackrel{{\scriptstyle(2)}}{{>}}\frac{(1-t)^{2}}{2}y$
@@
$\frac{\partial^{2}h(ty,y)}{\partial y^{2}}\geq$ 00 0
@@ Appendix C Properties of the Bellman Operator: KL-Case
    Targeted Value Function Analysis: Instead of considering the entire value function space $[0,(1-\gamma)^{-1}]^{\mathbf{S}}$, we restrict analysis to $V_{\mathcal{P}}^{\pi}$ specifically. This allows us to replace the $(1-\gamma)^{-1}$ dependence with the span semi-norm of $V_{P}^{\pi}$.
@@
$-\frac{p_{s,a}\left[\frac{V}{\alpha^{*}}e^{-V/\alpha^{*}}\right]}{p_{s,a}\left[e^{-V/\alpha^{*}}\right]}-\log p_{s,a}\left[e^{-V/\alpha^{*}}\right]=\delta$
@@ D.1 DR-DMDP under KL Uncertainty Set
Both $\text{Both errors}\leq O\left(\max_{\pi\in\Pi}\left\|V_{\widehat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}\right)$
@@ D.2 Reduction to DR-DMDP Approach under KL Uncertainty Set
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $96t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (D.6)
$\left\|\frac{V_{\widehat{\mathcal{P}}}^{*}}{\sqrt{n}}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq$ $48t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
00 0 $\leq g_{\mathcal{P}}^{\pi}(s)-(1-\gamma)V_{\mathcal{P}}^{\pi}(s)$ (D.14)
@@
00 0 $\geq g_{\mathcal{P}}^{\pi}(s)-(1-\gamma)V_{\mathcal{P}}^{\pi}(s)$ (D.15)
@@
$\leq$ $\frac{36t_{\mathrm{minorize}}}{\sqrt{n}}+36t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
$\leq$ $96t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $96t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (D.22)
$\left\|\frac{V_{\widehat{\mathcal{P}}}^{*}}{\sqrt{n}}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq$ $48t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@ D.3 Anchored DR-AMDP Approach under KL Uncertainty Set
As $\mathcal{P}$ is uniformly ergodic, for all $Q\in\mathcal{P}$, $Q$ is uniformly ergodic. Thus, for $\underline{Q}\in\underline{\mathcal{P}}$, by Lemma A.3, for any $\pi\in\Pi$ there exists an $(m_{\pi},p_{\pi})$ pair such that $\frac{m_{\pi}}{p_{\pi}}=t_{\mathrm{minorize}}(Q_{\pi})$. As for all $\underline{Q}\in\underline{\mathcal{P}}$, $\pi\in\Pi$, and $(s_{0},s_{m_{\pi}})\in\mathbf{S}\times\mathbf{S}$:
@@
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $96t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (D.31)
$\left\|g_{\underline{\widehat{\mathcal{P}}}}^{*}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq$ $48t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
Since for all $Q\in\mathcal{P}$, by Proposition B.1, we know $p_{s,a}\ll q_{s,a}$ for all $(s,a)\in\mathbf{S}\times\mathbf{A}$. Then by Lemma A.4, for all $Q\in\mathcal{P}$, $Q$ is also uniformly ergodic, and $\mathrm{Span}\left(g_{\mathcal{P}}^{\pi}\right)=0$. Thus:
@@
$\left\|g_{\underline{\widehat{\mathcal{P}}}}^{\pi}-g_{\mathcal{P}}^{\pi}\right\|_{\infty}\leq$ $\frac{18t_{\mathrm{minorize}}}{\sqrt{n}}+18t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (D.36)
@@
$\leq$ $24t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
$\leq$ $48t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
$\left\|g_{\underline{\widehat{\mathcal{P}}}}^{*}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq\max_{\pi\in\Pi}\left\|g_{\underline{\widehat{\mathcal{P}}}}^{\pi}-g_{\mathcal{P}}^{\pi}\right\|_{\infty}\leq 48t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (D.38)
@@ Appendix E Uniform Ergodicity of the $f_{k}$ Uncertainty Set
$\frac{dp_{k}(t)}{dt}\bigg{|}_{t=1}$ $=0\quad\text{ and }\quad\frac{dp_{k}(t)}{dt}\leq 0$ (E.4)
@@
When $21<k<$ 2, the second order derivative $d_{t}^{2}q_{k}(t)>0$ on $t\in(0,1)$. $dq_{k}(t)$ is monotone increasing, and as $d_{t}q_{k}(t)|_{t=1}=0$, we have $d_{t}q_{k}(t)\leq 0$ on $t\in(0,1]$, which sequentially implies $q_{k}(t)$ is monotone decreasing,
@@
And further $d_{t}h_{k}(t)\leq 0$, $h_{k}(t)$ is monotone decreasing from $\frac{1}{k}$ to $\frac{1}{2}$. Thus, for $21<k<$ 2:
@@ Appendix F Properties of the Bellman Operator: $f_{k}$-Case
00 0 $=\frac{\partial f(p_{s,a},V,\alpha)}{\partial\alpha}\bigg{|}_{\alpha=\alpha^{*}}$ (F.6)
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\alpha^{*}-c_{k}(\delta)p_{s,a}\left[(\alpha^{*}-V)_{+}^{k^{*}}\right]^{\frac{1}{k^{*}}}$
@@
Second, if $V$ is not a constnat, then $\mathbf{S}\backslash U\neq\emptyset$, observe that, by continuity, there exists an $\epsilon_{0}>0$, such that for all $0<\epsilon\leq\epsilon_{0}$,
@@
00 0 $\geq 1-c_{k}(\delta)\left(p_{s,a}\left[(\alpha^{*}+\varepsilon-V)_{+}^{k^{*}}\right]^{\frac{1}{k^{*}}-1}\cdot p_{s,a}\left[(\alpha^{*}+\varepsilon-V)_{+}^{k^{*}-1}\right]\right)$ (F.10)
@@
When $V$ is not essentially the constant, let $U=\left\{{s^{\prime}:V(s^{\prime})=\operatorname*{ess\,inf}_{p_{s,a}}V}\right\}$, and $\mathbf{S}\backslash U\neq\emptyset$, hence:
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sup_{\alpha\geq 2k^{*}\mathrm{Span}\left(V\right)}\frac{1}{k^{*}}\cdot\left|\frac{\Delta\left[\alpha^{k^{*}}\left(-k^{*}\frac{V^{\prime}}{\alpha}+\frac{k^{*}(k^{*}-1)}{2}(1-\xi)^{k^{*}-2}\frac{V^{\prime 2}}{\alpha^{2}}\right)\right]}{p\left[\frac{\alpha^{k^{*}-1}}{2}\right]}\right|$
@@
$G(t):=$ $\sup_{\alpha\geq\left\|V\right\|_{L^{\infty}(p_{s,a})}}g_{s,a}(t,\alpha)$ (F.21)
@@
$\leq$ $12t_{\mathrm{minorize}}(P_{\pi})\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$

### v2

@@ 2 Literature Review
We provide a summary of state-of-the-art sample complexity results in the literature in Table 1. In particular, we establish the first sample complexity guarantees for the average-reward DR-RL formulation, which achieves optimal dependence for $\varepsilon$ and $|\mathbf{S}||\mathbf{A}|$.
Table 1: Summary of S.O.T.A. sample complexity results in the literature, where $t_{\mathrm{mix}}$ is defined in  3.2.
@@
Standard
Discounted $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-3}\epsilon^{-2}})$ Azar et al., (2013); Li et al., (2020)
@@
DR-RL
Discounted $\widetilde{O}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-4}\epsilon^{-2}})$ Shi and Chi, (2024); Wang et al., 2024c
@@ 3.2 Uniform Ergodicity
An MDP (or its controlled transition kernel $P$) is said to be uniformly ergodic if for all policies $\pi\in\Pi$, $t_{\mathrm{minorize}}(P_{\pi})<\infty$. Then, define $t_{\mathrm{minorize}}:=\max_{\pi\in\Pi}t_{\mathrm{minorize}}(P_{\pi})<\infty$.
@@ 3.3 Distributionally Robust Discounted-Reward and Average-Reward MDPs
$\mathcal{T}_{\gamma}^{*}(V)(s)$= $\max_{a\in\mathbf{A}}\left\{{r(s,a)+\gamma\Gamma_{\mathcal{P}_{s,a}}(V)}\right\}$ (3.3)
$\mathcal{T}^{*}(v)(s)$= $\max_{a\in\mathbf{A}}\left\{{r(s,a)+\Gamma_{\mathcal{P}_{s,a}}(v)}\right\}$
@@
A DR-AMDP (or $\mathcal{P}$) is said to be uniformly ergodic if for all controlled kernels $\in\mathcal{P}, QQ$ is uniformly ergodic as in Definition 3.3.
@@ 4.2 Anchored DR-AMDP
for a solution pair $(g_{\underline{\widehat{\mathcal{P}}}}^{*},v_{\underline{\widehat{\mathcal{P}}}}^{*})$ .
@@
Sketched Proof of Theorems 4.4 and 4.5. Our proof begin with establishing that, under Assumption 2, each adversarial transition kernel $Q\in\mathcal{P}$ consists of conditional distributions $q_{s,a}$ that are absolutely continuous with respect to the nominal distributions $p_{s,a}$, with a uniform lower bound $1-\frac{1}{2m_{\vee}}$ on its Radon-Nikodym derivative. This guarantees that $t_{\mathrm{minorize}}(Q_{\pi})\leq O(t_{\mathrm{minorize}})$ for all $Q\in\mathcal{P}$ and $\pi\in\Pi$.
Next, combining Theorem D.1 with Lemma C.3, we establish that the policy error satisfies
$\left\|V_{\mathcal{P}}^{*}-V_{\mathcal{P}}^{\widehat{\pi}^{*}}\right\|_{\infty}\leq\frac{2}{1-\gamma}\max_{\pi\in\Pi}\left\|\widehat{\mathcal{T}}_{\gamma}^{\pi}(V_{\mathcal{P}}^{\pi})-\mathcal{T}_{\gamma}^{\pi}(V_{\mathcal{P}}^{\pi})\right\|_{\infty}.$
This reduces the analysis of the policy error to bounding the estimation error of the DR Bellman operator evaluated at $V_{\mathcal{P}}^{\pi}$. As the rewards cancel out, it remains to show that the DR functional applied to $V_{\mathcal{P}}^{\pi}$ satisfy appropriate concentration bound.
To this end, we apply the strong duality for the DR functional, the bound in Lemma C.6, and a Bernstein-type inequality to show that for any function $V$, the deviation satisfies
$\left\|\Gamma_{\widehat{\mathcal{P}}_{s,a}}(V)-\Gamma_{\mathcal{P}_{s,a}}(V)\right\|_{\infty}\leq\widetilde{O}\left(\frac{\mathrm{Span}\left(V\right)}{\sqrt{n\mathfrak{p}_{\wedge}}}\right)$
with high probability.
Finally, by selecting the parameters $\gamma=1-1/\sqrt{n}$ and $\eta=1/\sqrt{n}$, and noting that $\mathrm{Span}\left(V_{\mathcal{P}}^{\pi}\right)\leq O(t_{\mathrm{minorize}})$, we complete the proof for the KL-divergence case of Theorems 4.4 and 4.5. The argument under the $f_{k}$-divergence formulation proceeds in an analogous manner.
@@ 5 Numerical Experiments
This example clarifies our use of $m_{\vee}$ in Definition 3.3: while $m_{\pi}\equiv 1$ for all $\mathfrak{p}\in(0,1/2]$, the minorization time $t_{\mathrm{minorize}}$ is unbounded, approaching infinity as $\mathfrak{p}$ goes to 0.
@@
We then perform regression on data points on each MDP instance with the same parameter $\mathfrak{p}$. The plots demonstrate the error converging with rate $n^{-1/2}$, evidenced by the slope of 1/2-1/2 in Figure 2 on a log-log scale. We observe a remarkably low variance around the regression line of both algorithms, given that each data point is a single independent run of the corresponding algorithm.
@@
In addition to these experiment, we also perform a larger scale experiment to stress test our algorithm. Due to space limitations, the report is provided in Appendix H.
@@ 6 Conclusion and Future Work
Acknowledgement
N. Si gratefully acknowledges the support from the Hong Kong Research Grants Council [Theme-based Research Scheme T32-615/24-R].
@@ Appendix A Notations and Basic Properties
$\mathrm{Span}\left(u\right)$= $\max_{1\leq i\leq d}(u_{i})-\min_{1\leq i\leq d}(u_{i})$ (A.1)
@@ Appendix B Uniform Ergodicity of the KL Uncertainty Set
$\stackrel{{\scriptstyle(2)}}{{>}}\frac{(1-t)^{2}}{2}y$
@@
$\frac{\partial^{2}h(ty,y)}{\partial y^{2}}\geq$ $0$
@@ Appendix C Properties of the Bellman Operator: KL-Case
    Targeted Value Function Analysis: Instead of considering the entire value function space $0,(1-\gamma)^{-1}]^{\mathbf{S}}$, we restrict analysis to $V_{\mathcal{P}}^{\pi}$ specifically. This allows us to replace the $(1-\gamma)^{-1}$ dependence with the span semi-norm of $V_{P}^{\pi}$.
@@
$\frac{p_{s,a}\left[\frac{V}{\alpha^{*}}e^{-V/\alpha^{*}}\right]}{p_{s,a}\left[e^{-V/\alpha^{*}}\right]}-\log p_{s,a}\left[e^{-V/\alpha^{*}}\right]=\delta$
@@ D.1 DR-DMDP under KL Uncertainty Set
$\text{Both errors}\leq O\left(\max_{\pi\in\Pi}\left\|V_{\widehat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}\right)$
@@ D.2 Reduction to DR-DMDP Approach under KL Uncertainty Set
[…]
@@
$0$ $\leq g_{\mathcal{P}}^{\pi}(s)-(1-\gamma)V_{\mathcal{P}}^{\pi}(s)$ (D.14)
@@
$0$ $\geq g_{\mathcal{P}}^{\pi}(s)-(1-\gamma)V_{\mathcal{P}}^{\pi}(s)$ (D.15)
@@
[…]
@@
[…]
@@
[…]
@@ D.3 Anchored DR-AMDP Approach under KL Uncertainty Set
As $\mathcal{P}$ is uniformly ergodic, for all $\in\mathcal{P}, QQ$ is uniformly ergodic. Thus, for $\underline{Q}\in\underline{\mathcal{P}}$, by Lemma A.3, for any $\pi\in\Pi$ there exists an $(m_{\pi},p_{\pi})$ pair such that $\frac{m_{\pi}}{p_{\pi}}=t_{\mathrm{minorize}}(Q_{\pi})$. As for all $\underline{Q}\in\underline{\mathcal{P}}$, $\pi\in\Pi$, and $(s_{0},s_{m_{\pi}})\in\mathbf{S}\times\mathbf{S}$:
@@
[…]
@@
Since for all $Q\in\mathcal{P}$, by Proposition B.1, we know $p_{s,a}\ll q_{s,a}$ for all $(s,a)\in\mathbf{S}\times\mathbf{A}$. Then by Lemma A.4, for all $\in\mathcal{P}, QQ$ is also uniformly ergodic, and $\mathrm{Span}\left(g_{\mathcal{P}}^{\pi}\right)=0$. Thus:
@@
[…]
@@
[…]
@@
[…]
@@ Appendix E Uniform Ergodicity of the $f_{k}$ Uncertainty Set
$\frac{dp_{k}(t)}{dt}\bigg|_{t=1}$ =0 $=0\quad\text{ and }\quad\frac{dp_{k}(t)}{dt}\leq$ 0 (E.4)
@@
When $1<k<2$, the second order derivative $d_{t}^{2}q_{k}(t)>0$ on $t\in(0,1)$. $dq_{k}(t)$ is monotone increasing, and as $d_{t}q_{k}(t)|_{t=1}=0$, we have $d_{t}q_{k}(t)\leq 0$ on $t\in(0,1]$, which sequentially implies $q_{k}(t)$ is monotone decreasing,
@@
And further $d_{t}h_{k}(t)\leq 0$, $h_{k}(t)$ is monotone decreasing from $\frac{1}{k}$ to $\frac{1}{2}$. Thus, for $1<k<2$:
@@ Appendix F Properties of the Bellman Operator: $f_{k}$-Case
$0$ $=\frac{\partial f(p_{s,a},V,\alpha)}{\partial\alpha}\bigg|_{\alpha=\alpha^{*}}$ (F.6)
@@
$\stackrel{{\scriptstyle(i)}}{{=}}\alpha^{*}-c_{k}(\delta)p_{s,a}\left[(\alpha^{*}-V)_{+}^{k^{*}}\right]^{\frac{1}{k^{*}}}$
@@
[…]
@@
$0$ $\geq 1-c_{k}(\delta)\left(p_{s,a}\left[(\alpha^{*}+\varepsilon-V)_{+}^{k^{*}}\right]^{\frac{1}{k^{*}}-1}\cdot p_{s,a}\left[(\alpha^{*}+\varepsilon-V)_{+}^{k^{*}-1}\right]\right)$ (F.10)
@@
[…]
@@
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sup_{\alpha\geq 2k^{*}\mathrm{Span}\left(V\right)}\frac{1}{k^{*}}\cdot\left|\frac{\Delta\left[\alpha^{k^{*}}\left(-k^{*}\frac{V^{\prime}}{\alpha}+\frac{k^{*}(k^{*}-1)}{2}(1-\xi)^{k^{*}-2}\frac{V^{\prime 2}}{\alpha^{2}}\right)\right]}{p\left[\frac{\alpha^{k^{*}-1}}{2}\right]}\right|$
@@
$G(t)$= $\sup_{\alpha\geq\left\|V\right\|_{L^{\infty}(p_{s,a})}}g_{s,a}(t,\alpha)$ (F.21)
@@
[…]